Calculate the surface area and volume of a rectangular solid measuring 43mm in length, 12mm in width, and 19mm in height (report your answer in cm square and cm cube). The mass of this block is 8.4g. What is it's density and will it sink or float in water?
Answer #1

Length of rectangular solid = 4.3cm

Width of rectangular solid = 1.2cm

Height of rectangular solid = 1.9cm

Surface area =2(wl+hl+hw) = 2( 1.2x4.3+1.9x4.3+1.9x1.2)cm2

                                             = 2(5.16+8.17+2.28)cm2

                                                 = 2x 15.61cm2

                                             = 31.22cm2

Volume = whl = 1.2x1.9x4.3cm3 = 9.804cm3

Density = Mass/Volume = 8.4g/9.804mm3 = 0.856g/cm3.

As the density of this solid is less than the density of water, so it will float.